# Regular triangle unions with maximal number of sides

### v3

@@
Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean (IDE) languages. During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration, where Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which has common IDE roots with English East. was shown. It forms an octagonal star (Fig. 1, left), which is just a standard octagram ($\{8,3\}$-gon in Schläfli notation). Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, of course, in magic; same picture, left, where simple and double Lietuvēns crosses are shown), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon. With few more thoughts, it is immediate how one can increase edge count by slightly deforming all triangles. It was therefore natural to wonder what is the best possible result. Such is a subjective genesis of the problem. And magic surely works! The author could not anticipate that hypostases of both crosses will eventually emerge in this purely mathematical endeavour. Yet, the earliest to rise will be $\ldots$ the Star of David. But let us go to the basics first.
@@
Take $n$ $\triangle$. Its edges will henceforth be called legs, while edge will be reserved solely for a $\bigcup$ polygon. Any two of $\triangle$ intersect in at most $6$ points, and it is straightforward to contrive an example with $\sim 2n^{2}$ edges on the boundary of the $\bigcup$. However, in such construction many of the $\triangle$ appear to be very slim. If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ then has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat . What happens, though, if the $\bigcup$ has no holes at all and its interior is connected? Let us formulate this precisely.
Problem.
@@
  • 4)
    Vertices are on the unit circle and are arranged in the cyclic way – seq. $r(n)$.
It may seem surprising that the question “how many edges exactly?" has not been asked before. Yet, there are reasons for that. The seq. $i(n)$ is the one where existing techniques and results apply most. Take a $\bigcup$ of $n$ $\triangle$ with $i(n)$ edges. Find a $\triangle$ which contributes least to the common boundary. Remove it. It is not difficult to see this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$. Thus, as if
@@
This would imply $i(n)=O(n\log n)$. Sure, such a naïve approach is not strictly valid: we do not know whether removal of 11 $\triangle$ leaves no holes in the $\bigcup$. But a certain aspect of the phenomenon is captured. In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called Ackermann function. It tends to $\infty$ with an argument, but extremely slowly. Its value is $3$ in the interval $[5,16]$, and $4$ in the interval
@@
It is impossible for a sane mind to fathom a place on the line of positive integers where this function surpasses $5$. There is a handful of other reasons why a plain answer is not to be expected.
@@
However, the situation with cyclically arranged vertices is slightly different. Such $\triangle$ $\bigcup$ will be called regular. The current paper, though mostly expository and derivative, introduces such objects for the first tiome and gives a new additional tool in dealing with $\bigcup$. Namely, triangulation shifts. Theoretically, an answer becomes within a reach. Sure, there exists problems in combinatorial geometry where an arm-stretch distance is still a considerable one.
For example, it is known that two simple polygons with odd numbers ($m$ and $n$) of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig. 2), and this is conjectured to be the peak value. Ackerman et al. ackerman proved that the maximum is indeed mnmn+Cmn-m-n+C. Accordingly, this is optimal, apart from the exact value of $C$ (still not rigorously known to be equal to $3$).
Henceforth, we will concentrate on $e(n)$, referred here as the Politrigon seq.. It now has a number A375986 in the OEIS oeis . What are its initial values? For $n=1$, there is nothing to prove. The $\bigcup$ of 22 $\triangle$, if a simple $M$-gon, can have any number of sides in the range $3\leq M\leq 10$, and also $M=12$. The latter is mostly familiar from a pleasingly symmetric hexagram known as the star of David. In order to prove that $e(3)=22$, case-by-case analysis can be performed. The example ($n=1,2,3$) is presented in Fig. 11 (three brown $\triangle$).
Figure 3: Regular $\bigcup$ of $4$ triangles with $33$ sides.
@@
Figure 5: All 44 perfect $\bigcup$ of triangles. These are exactly the configurations satisfying (n+1)e(n)=(12n6)n(n+1)e(n)=(12n-6)n. Allegorically speaking, the line of succession to the Eye of Providence continues with the Star of David, then with the Pentastar, and terminates with the Octastar. The last two have only rotational symmetry. Combinatorics of Pentastar cannot be reached from Octastar by removing 33 $\triangle$. This excludes any inductive construction of the best $\bigcup$.
Moving onto $n=4$, similar computer-assisted overhaul show that $e(4)=33$ (Fig. 3). If we knew all possible arrangements of $3n$ lines in the affine plane (seq. A090338), this catalogue would encode, among other things, the Politrigon seq. (and $i(n)$, too). Such list (for $n$ lines) would also encode the solution to the Kobon $\triangle$ problem (seq. A006066). The latter asks how many non-overlapping $\triangle$ can be formed by $n$ lines (see Fig. 4 depicting one of maximal configurations for $n=10$). In its full generality, Kobon’s problem is still unsolved. It is obvious that such full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number. In fact, the last exact value known is for $9$ lines, which amounts to $3.111.341$ simple arrangements. In our setting, this would settle, with some effort, the case $n=3$. Clearly, in all derivatives of line arrangement problem, other method should be looked for.
And it does exist in regular $\triangle\bigcup$ case. With such tools at hand, it will be possible to demonstrate the unconditional inequality $(n+1)e(n)\leq(12n-6)n$. This leads to the discovery of two “perfect" $\triangle$ $\bigcup$ showing that $e(5)=45$ and $e(8)=80$ are possible. Together with a $\triangle$ and a hexagram, this is the complete list of maximally saturated $\triangle$ $\bigcup$. In other words, the only cases where the equality is achieved (Fig. 5). (exactly the very reincarnation of two magic crosses hinted to in the Introduction). For $n\neq 1,2,5,8$, strict inequality must hold.
@@
The seq. $e(n)$ starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$. It also satisfies the following:
$11\leq e(n+1)-e(n)\leq 14\quad(forn\geq 3),$
$e(n)\leq 12n-18+I(10\leq n\leq 13)\quad(forn\geq 10).$
The first inequality gives situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$). This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
The upper bound now stands in sharp contrast with a superlinear growth of the function $i(n)$. It not just bounds $e(n)$ from above, but in fact gives the exact combinatorial bound on first approximation. Without explaining this vague phrase now, here is a little trick to be played on the readers. Among all triangulations of a regular $n$-gon, choose the one whose maximal sum of number of diagonals adjacent at three consecutive vertices is minimal. Denote this minimum by $\delta(n)$ (see Fig. 6). One of the many interpretations of Catalan numbers $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ is that $C_{n-2}$ gives the number of triangulations of a convex $n$-gon. Since there are n3n-3 diagonals, $\delta(n)\geq\big\lceil 6-\frac{18}{n}\big\rceil$. The following statement is therefore a result of full check ($n\leq 17$) and explicit construction ($n\geq 18$):
Proposition 1.
The sequence $\delta(n)$, $n\geq 3$, starts as $0,2,3,3,4,5,4,5,5,5,5,5$, and then stabilizes at the value $6$ (see Fig. 6).
The reader can verify that for $n\geq 2$, $\mathscr{U}(n)=12n-12-\delta(n+1)$ coincides with exact value of the seq. $e(n)$ and gives the exact upper combinatorial bound for $n\geq 10$. The relation of $\triangle$ $\bigcup$ to triangulations of $n$-gons is unexpected, and will be clarified later. But now, let us discuss what does it mean “combinatorial" vs. “geometric". [^3] Since we started from linguistics, it is apt to emphasize here that suffix dilemma ‘-ial’ vs. ‘-ic’ essentially coincides with whether it is of Latin or of Greek origin. In other words, what are pseudolines and what is stretchability.
Figure 6: Triangulations of regular polygons which minimize the maximum of number of diagonals adjacent at three consecutive vertices. This minimum is denoted by $\alpha(n)$
@@
Apart from A006066 and A090338, there are other seq. in the OEIS made from kindred ingredients. This means: seq., whose definition employs only well-familiar objects from Euclidean geometry, but whose calculation requisites rather delicate arrangements of these. One of such seq. was popularized by Neil Sloane and the community of OEIS during the “ranking process" invoked to clarify which seq. should be assigned a honorary label A250.000. A runner-up in this vote, now a seq. A250.001 (first proposed by Jonathan Wild), asks to find the number of ways to draw $n$ circles in the affine plane sloane ; oeis (definitions of admissible and equivalent configurations are intuitive and can be found in the references cited). In Figure 7, three (out of $173$) arrangements of $4$ circles are reproduced.
@@
Even having a ready example in front of the eyes, it takes some practice to correctly drag around centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$. See a recent survey cutler for many more examples of such seq., associated with cutting of plane with all kinds of “exotic knives".
On the other hand, pseudoline arrangements and stretchability problem are not only close in spirit, but play a significant role in determining $e(n)$. A pseudoline is a simple connected curve in the projective plane such that its removal does not disconnect $\mathbb{R}P^{2}$. Arrangement of pseudolines is a finite set of them so that each two meet exactly once, and there are at least two such intersection points (the affine case is modified accordingly). A survey by Felsner and Goodman felsner is another enriching reading. The most transparent arrangement of projective pseudolines that cannot be realized with lines (that is, “non-stretchable") is given by a twisted Pappus configuration as in Fig. 8 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of a recent Intelligencer schwartz ).
@@
Figure 9: Non-stretchable simple arrangement of $9$ affine pseudolines
Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed). If $n\leq 7$, any one of them can be realized with lines. Yet, the configuration in Fig. 9 (found by L. Reeves and J. Wild again) with $8$ pseudolines cannot be straightened. A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant dallant and is $2^{0.6496n^{2}}$. It is based on Zone Theorem (that will be discussed later). Compare this bound to a one given earlier for simple arrangements of $n$ lines.
The reason for introducing pseudolines is that their arrangement problems are easier and some aspects can be tackled solely combinatorially. we will witness this. It appears that many (but not all!) results about lines are more about their configuration that straightness property itself. There are other reasons to introduce pseudolines. For example, Suvorov suvorov gives two simple arrangements of $13$ lines in the affine plane. As pseudolines, one can be transformed into the other via a continuous deformation of the plane, so that the configuration always remains simple (flexible isotopy). This, however, cannot be done if curves are always required to remain straight (rigid isotopy).
Take the $\bigcup$ of $n$ $\triangle$ labelled $1$ through $n$. Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the next edge belongs. The result of this paper amounts to constructing a seq. of maximally possible length $\mathscr{U}(n)$ which serves as a model of such actual seq. The second step would be to show that this model is realisable in pseudoline geometry. Possibly, this is not too hard to show. The question whether a construction can further be straightened is of a very different sort. Stretchability problem is known to be NP-hard.
@@
It is straightforward to contrive $n\geq 1$ $\triangle$ whose $\bigcup$ is a simple polygon with 9n69n-6 edges. One only needs to choose $3$ points on $3$ different sides of the existing polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig. 10A). The only two concerns are that no vertex of an initial polygon lie inside the newly constructed $\triangle$, and that the new polygon is simple. If one vertex of the initial polygon with reflex ($180^{\circ}$) angle lies inside a $\triangle$, we get an increase of edge count by $10$ (Fig. 10B). If two non-consecutive vertices with reflex angles are inside, the increase is by $11$ (Fig. 10C).
@@
With this idea in mind, let us take, as a start-up, three brown $\triangle$ as shown in Fig. 11. Their $\bigcup$ has $22$ edges. Consider a blue segment and two green circles as shown. The latter are inscribed into $\angle PXA$ and $\angle SYB$, each passing through points $A$ and $B$, respectively. Next, choose several points (labelled $1,2,\ldots,m=n-3$, left to right) strictly inside the blue segment. From every point $p$, draw “inner" tangents to both circles. Let the intersection with $PQ$ and $SR$ be, respectively, $p^{\prime}$ and $p^{\prime\prime}$. Fig. 11 shows example with $m=2$.
@@
We now miniaturelly protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
In fact, nothing is specific about this start-up configuration. The construction works each time there exists an edge which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$). Take, for example, the green edge in Fig. 12 (the one with a control vertex numbered “47"). Its endpoints belong to a cyan and grey $\triangle$, respectively. Take a point $x$ inside this edge. We will construct a $\triangle$ whose vertex is $x$, only slightly protruded outwards. It will lie between black and brown vertices $a$ and $b$, respectively. Let left and right edges at these vertices intersect at point $p$ and $q$ (shown also on the same Fig. 12). Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold) and black segments on the boundary of the union. Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count. This mimics precisely the construction in Fig. 11. Also, vertices of triangles preserve a cyclic order.
@@
Note that such lower bounds are situational. To see this, consider an Octastar in Fig. 5. Removing any 33 $\triangle$ give a collection of 55 $\triangle$ with at most $44$ edges in the $\bigcup$. In other words, we cannot expect to achieve maximum number of edges for $(n+1)$ $\triangle$ by finding the best addition to the existing best $n$ $\triangle$ configuration. Each $n$ should be treated individually, what we will do in the section ‘Maximal Sequences’.
@@
Take a regular $\bigcup$ of $n$ $\triangle$ with $M$ edges. For example, consider 88 $\triangle$ in Fig. 12. We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with control label “10". All $\triangle$ are numbered $1$ through $8$. So, for example, black $\triangle$ gets number $4$. Now, go along the boundary in a positive direction and record what edges occur. This produces the cyclic seq. of length $79$:
@@
$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}.$ (1)
@@
Take a fragment of the boundary contained between two vertices of the same blue $\triangle$. Say, the simply-underlined one in (1). Write down numbers of $\triangle$ whose RIGHT edges are encountered along the way. That is: ignore a number if its double has not yet occurred; otherwise, include it; from each doubled number, only one copy is included. This way the seq. $\mathcal{L}=123214546417871$ is obtained. What properties are intrinsic to it? First, no two adjacent numbers are equal. Second, though some of them may (and do) appear multiple times, such seq. do not contain a subseq. of the form abababab for any distinct pair $a,b\in\{1,2,\ldots,n\}$. This is a crucial property of regular $\triangle$ $\bigcup$. In the section ‘Lower envelopes’ we will see that as long as such seq. are allowed to have subseq. of the form abababab, their length can grow superlinearly.

### v4

@@
Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean languages. During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration depicting Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which is cognate with English East.. It forms an octagonal star (Fig. 1, left), which is just a standard octagram ($\{8/3\}$-gon in Schläfli notation). Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, undoubtedly, in magic; same picture, left, where simple and double ‘Lietuvēns’ crosses are shown [^2] For a background, see English Wikipedia page with this title.), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon. With few more thoughts, it is immediate how to increase an edge count by slightly deforming all triangles. It was therefore natural to wonder what is the best possible result. Such is a subjective genesis of the problem. And the magic surely works! The author could not anticipate that hypostases of both crosses will eventually emerge in this purely mathematical endeavour. Yet, the earliest to rise will be $\ldots$ the Star of David. But let us go to the basics first.
@@
Take $n$ $\triangle$. Its sides will henceforth be called legs, while edge will be solely reserved solely for a $\bigcup$. Any two of $\triangle$ intersect in at most $6$ points, and it is straightforward to contrive an example with $\sim 2n^{2}$ edges on the boundary of the $\bigcup$. However, in such construction many of the $\triangle$ appear to be very slim. If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat . What happens, though, if the $\bigcup$ has no holes at all and its interior is connected? Let us formulate this precisely.
Problems..
@@
  • $\mathrm{4)}$
    Vertices are on the unit circle and are arranged in a cyclic way – seq. $r(n)$.
It may seem surprising that the question “how many edges exactly?" has not been asked before. Yet, there are reasons for that. The seq. $i(n)$ is the one where existing techniques and results apply most. Take a $\bigcup$ of $n$ $\triangle$ with $i(n)$ edges. Find a $\triangle$ which contributes least to the common boundary. Remove it. It is not difficult to see that this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$. Thus, as if
@@
This would imply $i(n)=O(n\log n)$. Sure, such a naïve approach is not strictly valid: we do not know whether removal of 11 $\triangle$ leaves no holes in the $\bigcup$. But a certain aspect of the phenomenon is captured. In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir & Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called (William) Ackermann function. There are several variants of the latter, and we do not give the precise definition. The one used here can be found in sharir , p. 12-15. It is not essential which variant of the Ackermann function is meant, all asymptotic results hold. The function $\alpha$ tends to $\infty$ with an argument, but extremely slowly. Its value is $3$ in the interval $[5,16]$, and $4$ in the interval
@@
It is impossible for a sane mind to fathom a place on the positive half-line where this function surpasses $5$. There is a handful of other reasons why a plain answer about $i(n)$ is not to be expected. It is even worse with the seq. $c(n)$. Apart from $c(n)\leq i(n)$, we do not know where the things stand. No $\Omega$ result is currently available.
@@
However, the situation with cyclically arranged vertices is different. Such $\triangle$ $\bigcup$ will be called regular. By this it is also implied that the $\bigcup$ is proper – all vertices are among vertices of the $\bigcup$ polygon. In contrast from the situation with $i(n)$, regular $\bigcup$ is automatically a simple polygon. Henceforth, $e(n)$ will be our main focus. It now has a number A375986 in the OEIS oeis . The current paper, though mostly expository, introduces regular $\bigcup$ for the first time and gives a new tool to particularly deal with it. Theoretically at least, the precise answer for any $n$ becomes within a reach.
Sure, there are problems in combinatorial geometry where an arm’s length distance is still a considerable one. For example, it is known that simple $m$-gon and $n$-gon with odd number of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig. 2), and this is conjectured to be the peak value. Eyal Ackerman et al. ackerman proved that the maximum is indeed mnmn+Cmn-m-n+C. Accordingly, this is optimal, apart from the exact value of $C$, still not rigorously known to be equal to $3$.
What are the initial values of $e(n)$? For $n=1$, there is nothing to prove. The $\bigcup$ of 22 $\triangle$, if a simple $M$-gon, can have any number of sides in the range $3\leq M\leq 10$, and also $M=12$. Only in the latter case the $\bigcup$ is regular, which is mostly familiar from a pleasingly symmetric hexagram known as the star of David. In order to prove that $e(3)=22$, case-by-case analysis can be performed. The example is presented in Fig. 10 (three brown $\triangle$).
Figure 3: Regular $\bigcup$ of 44 triangles with $33$ sides. This unique $33$-gon is calibrated to maximize the ratio of its area to a sum of squares of its edge-lengths.
@@
Figure 5: All 33 perfect proper unions of triangles. These (along with a single $\triangle$) are exactly the unions satisfying (n+1)M=(12n6)n(n+1)M=(12n-6)n. The pentagram reincarnates as the Pentastar, and the Octastar is an unexpected hypostasis of Auseklis.
Moving onto $n=4$, similar computer-assisted overhaul shows that $e(4)=33$ (Fig. 3). If we knew all possible arrangements of $3n$ lines in general position in the affine plane (seq. A090338), this catalogue would encode, among other things, the seq. $e(n)$ (and $i(n)$, too). Such list (for $n$ lines) would also encode the solution to the Kobon $\triangle$ problem (seq. A006066). The latter asks how many non-overlapping $\triangle$ can be formed by $n$ lines (see Fig. 4 depicting one of maximal configurations for $n=10$). In its full generality, Kobon’s problem is still unsolved. It is obvious that such a full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number. In fact, the last exact value known is for $9$ lines, which amounts to $3.111.341$ simple arrangements. In our setting, this would settle, with some effort, the case $n=3$. Clearly, in any derivative of the line arrangement problem, other method should be looked for. And here is the main tool.
Definition..
Let $n\geq 2$. Given a (possibly, incomplete) triangulation $\mathcal{T}_{0}$ of a regular $(n+1)$-gon and a distinguished vertex $1$, other vertices labelled $2$ to $n+1$ in a positive direction. A $(n+1)$-triangulation shift is a collection of $3n+1$ triangulations $\mathcal{T}_{i}$ ($0\leq i\leq 3n$) on the same set of vertices. $\mathcal{T}_{i}$ is obtained from $\mathcal{T}_{i-1}$ by removal and a subsequent addition of some diagonals incident to a vertex $i$. The final triangulation $T_{3n}$ must coincide with $T_{0}$, but turned $3$ vertices clockwise. If $t_{i}$ is the number of diagonals of $\mathcal{T}_{i}$ incident to $i$, the weight of such a shift is defined to be $\omega(\mathcal{T}_{0})=\sum_{i=1}^{3n}t_{i}+6n$.
So, we go round a regular $(n+1)$-gon exactly $3$ vertices short of $3$ full turns. It is allowed to slightly alter diagonals at any present vertex. The final result is the initial triangulation, only rotated by 3-3 turns. Intuitively, the biggest weight is achieved for those $\mathcal{T}_{0}$ which best coincide with itself after such a rotation. This object is of interest, but seems coming out of the blue. What’s its relation to $\triangle$ $\bigcup$? The answer will gradually unfold.
@@
The seq. $e(n)$, $n\geq 1$, starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$. Given any regular $\triangle$ $\bigcup$ with $M$ edges, there exists a $(n+1)$-triangulation shift with weight $M$. For $n\geq 5$,
$e(n)\leq 11n-14+2\bigg\lfloor\frac{n+1}{3}\bigg\rfloor=\mathscr{U}(n),$
and $\mathscr{U}(n)$ is the maximal weight among all $(n+1)$-triangulation shifts. Additionally, the inequality $11\leq e(n+1)-e(n)\leq 14$ holds.
The last stated result gives a situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$). This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
In the course of the proof, we will frequently aberrate to cognate topics. Yet, the reader can be rest assured that there is no complicated scheme behind the proof. All steps will align into a transparent row. As an intermediate result, we will show that $(n+1)e(n)\leq(12n-6)n$. This a posteriori leads to the discovery of two “perfect" $\triangle$ $\bigcup$ and a demonstration that $e(5)=45$, $e(8)=80$. Together with a $\triangle$ and a hexagram, the four form the complete list of maximally saturated (that is, where the equality is achieved) $\triangle$ $\bigcup$ ( Fig. 5). This exactly corresponds to $(n+1)$-triangulation shifts where no diagonal is removed or added! In other words, to a triangulation which, if rotated by 3-3 vertices, coincides with itself. For $n\neq 1,2,5,8$, strict inequality must hold.
The asserted upper bound stands in sharp contrast with a superlinear growth of the function $i(n)$. The function $\mathscr{U}(n)$ can be thought as being the exact combinatorial bound on the first approximation. Let us now discuss what does it mean “combinatorial" vs. “geometric". [^4] Quite a symbolic collision. On a linguistic level, the clash of suffixes ‘-ial’ vs. ‘-ic’ is essentially equivalent to whether the root is of Latin or of Greek origin. In other words, what are pseudolines and what is stretchability.
@@
Apart from A006066 and A090338, there are other seq. in the OEIS made from kindred ingredients and emitting alike aroma. This means: seq., whose definition employs only well-familiar objects from Euclidean geometry, but whose calculation requisites rather delicate arrangements of these. One of such seq. was popularized by Neil Sloane and the community of OEIS during the “ranking process" invoked to clarify which seq. should be assigned a honorary label A250.000. A runner-up in this vote, now a seq. A250.001 (first proposed by Jonathan Wild), asks to find the number of ways to draw $n$ circles in the affine plane sloane ; oeis (definitions of admissible and equivalent configurations are intuitive and can be found in the references cited). In Fig. 6, three (out of $173$) arrangements of $4$ circles are reproduced.
@@
Even having a ready example in front of the eyes, it takes some practice to correctly drag with mouse centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$. See a recent survey cutler for many more examples of such seq., associated with cutting of plane with all kinds of “exotic knives".
Going further, problems like pseudoline arrangements and stretchability are not only close in spirit and delicacy level, but play a significant role in determining $e(n)$. A pseudoline is a simple connected curve in the projective plane such that its removal does not disconnect $\mathbb{R}P^{2}$. Arrangement of pseudolines is their finite set so that each two meet exactly once, and it is not a pencil (so, there must be at least two intersection points). The affine (Euclidean) case is modified accordingly. A survey by Felsner and Goodman felsner is another enriching reading. The most transparent arrangement of projective pseudolines that cannot be realized with lines (that is, “non-stretchable") is given by a twisted Pappus configuration as in Fig. 7 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of Intelligencer schwartz ).
@@
Figure 8: One of 33 (out of $38612$) non-stretchable simple arrangement of $8$ affine pseudolines
Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed; seq. A090339). [^5] There are subtlenesses in what is being counted by any particular seq. If $n$ affine lines are in general position, there are no finite multiple intersection points, apart from that some subsets may be parallel. So, if treated as projective lines, such arrangement might not be simple. There are sequences which count arrangements with a cell distinguished and (un)oriented, etc. If $n\leq 7$, any one of them can be realized with lines. Yet, the configuration in Fig. 8 (found by L. Reeves and J. Wild) with $8$ pseudolines cannot be straightened. A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant and is $2^{0.6496n^{2}}$ dallant . The proof is based on the Zone Theorem (that will be discussed later). Compare this bound to a one given earlier for simple arrangements of $n$ straight lines (as we will see, in the light of this discrepancy, the aforementioned result of Wiernik and Sharir becomes even more surprising).
The reason for introducing pseudolines is that their arrangement problems are easier and some aspects can be tackled solely combinatorially. It appears that many (but not all!) results about lines are more about their configuration that straightness itself. There are other reasons to introduce pseudolines. For example, Suvorov suvorov gives two simple arrangements of $13$ lines in the affine plane. As pseudolines, one can be transformed into the other via a continuous deformation of the plane, so that the configuration always remains simple (flexible isotopy). This, however, cannot be done if curves are always required to remain straight (rigid isotopy). See the end of Chapter 4 in felsner for few more spotlight distinctions.
Take the $\bigcup$ of $n$ $\triangle$ labelled $1$ through $n$. Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the consequent edge belongs. The result of this paper amounts to constructing a seq. of maximal possible length $\mathscr{U}(n)$ which serves as a model of such actual seq. The second step would be to show that this model is realisable in pseudoline geometry. This might be not too hard to show. The question whether a construction can further be straightened is of a very different sort. The Section ‘Lower envelopes’ will contain an example of such a problem. Stretchability problem is known to be NP-hard.
@@
It is straightforward to contrive $n\geq 1$ $\triangle$ whose $\bigcup$ is a simple polygon with 9n69n-6 edges. One only needs to choose $3$ points on $3$ different sides of the existing simple polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig. 9A).
@@
The only two concerns are that no vertex of an initial polygon lie inside the newly constructed $\triangle$, and that the new polygon is simple. If, however, two incident edges are intersected by the same leg of a new $\triangle$, we get an increase of edge count by $10$ (Fig. 9B). If two non-overlapping pairs of incident edges are intersected, the increase is by $11$ (Fig. 9C).
With this idea in mind, let us take, as a start-up, three brown $\triangle$ as shown in Fig. 10. Their $\bigcup$ has $22$ edges. Consider a blue segment and two green circles as shown. The latter are inscribed into $\angle PXA$ and $\angle SYB$, each passing through points $A$ and $B$, respectively. Next, choose several points (labelled $1,2,\ldots,m=n-3$, left to right) strictly inside the blue segment. From every such point $p$, draw inner tangents to both circles. Let the intersection with $PQ$ and $SR$ be, respectively, $p^{\prime}$ and $p^{\prime\prime}$. Fig. 10 shows example with $m=2$.
@@
We now minutely protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
In fact, nothing is specific about this start-up configuration. The construction works each time there exists an edge of the $\bigcup$ which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$). Take, for example, the green edge in Fig. 11 (the one with a control vertex numbered “47"). Its endpoints belong to a cyan and grey $\triangle$, respectively. Take a point $x$ inside this edge. We will construct a $\triangle$ whose vertex is $x$, only slightly protruded outwards. It will lie between black and brown vertices $a$ and $b$, respectively. Let left and right edges at these vertices intersect at point $p$ and $q$ (shown also on the same Fig. 11). Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold just a bit to the left of $p$) and black edges of the $\bigcup$. Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count, also vertices of $\triangle$ preserve a cyclic order. This mimics precisely the construction in Fig. 10.
@@
Take a regular $\bigcup$ of $n$ $\triangle$ with $M$ edges. For example, consider 88 $\triangle$ in Fig. 11. We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with a control label “10". All $\triangle$ are numbered $1$ through $8$. For instance, black $\triangle$ gets number $4$. Now, go along the boundary in a positive direction and sequentially record what edge occurs. This produces the cyclic seq. of length $79$:
@@
$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}$ (1)
@@
Take separately a fragment of this boundary contained between two vertices of the same blue $\triangle$. Say, the singly-underlined one in (1). Write down numbers of $\triangle$ whose RIGHT edges are encountered along the way. That is: ignore a number if its double has not yet occurred; otherwise, record it; from each doubled number, only one copy is included. This way the seq. $\mathcal{L}=123214546417871$ is obtained. What properties are intrinsic to it? First, no two adjacent numbers are equal. Second, though some numbers may (and do) appear multiple times, such seq. do not contain a subseq. of the form abababab for any distinct pair $a,b\in\{1,2,\ldots,n\}$. This is a crucial attribute of a regular $\triangle$ $\bigcup$. In the section ‘Lower envelopes’ we will see that as long as such seq. are allowed to have subseq. of the form abababab, their length can grow superlinearly.